OpenStudy (anonymous):

Solve the following for x: 3 – (x +(-4)) = 2x – 6 + 5x Explain in as much detail as possible.

OpenStudy (across):

You have:\[3-(x+(-4))=2x-6+5x.\]Simplify the left hand side by expanding the parentheses:\[3-x-(-4)=2x-6+5x,\\3-x+4=2x-6+5x.\]Combine like terms in both the left and right hand sides:\[7-x=7x-6.\]Solve for \(x\):\[13=8x,\\x=?\]